OpenStudy (anonymous):

The ordered pairs below represent a relation between x and y.(-3,-3), (-2,0), (-1,3), (0,6), (1,9), (2,12)Could this set of ordered pairs have been generated by a linear function?

OpenStudy (mertsj):

you could choose two of the points and write the equation of the line joining those points. Then plug in the other points and see if they make the sentence true.

OpenStudy (mertsj):

Let's choose (1,9) and (2,12) The slope if 12-9/2-1 or 3

OpenStudy (mertsj):

y-9=3(x-1) y=3x+6

OpenStudy (mertsj):

Now plug in the other points and see if they are on that line.

OpenStudy (anonymous):

can you just give me the final answer & why?

OpenStudy (mertsj):

did you plug in the other points?

OpenStudy (mertsj):

They all work. so yes, those points are collinear.

OpenStudy (mertsj):

Because they all make the sentence y=3x+6 true